OpenStudy (anonymous):

Find the product of the complex number and its conjugate. 1 + 3i

OpenStudy (anonymous):

the conjugate of \(a+bi\) is \(a-bi\) and it is always true that \[(a+bi)(a-bi)=a^2+b^2\]

OpenStudy (anonymous):

forget about the \(i\) and forget about minus signs, in your case \(a=1,b=3\) and \(1^2+3^2=1+9=10\)

OpenStudy (anonymous):

So is it 1+9i?

OpenStudy (jim766):

right..

OpenStudy (anonymous):

That's the answer? Your ".." is misleading.

OpenStudy (anonymous):

there is no "\(i\)" in the answer

OpenStudy (anonymous):

no \(i\) and no \(-\) sign, only \(a^2+b^2\)

OpenStudy (jim766):

sorry my bad.... i^ 2 = 1

OpenStudy (anonymous):

\((a+bi)(a-bi)=a^2+b^2\) so \((1+3i)(1-3i)=1^2+3^2=1+9=10\) a real number

OpenStudy (jim766):

1 + 9 = 10

OpenStudy (anonymous):

1 + 9i 10 -8 1 - 9i Are my answers, is it 10?

OpenStudy (anonymous):

yes, it is 10

OpenStudy (anonymous):

Thank you!